PREDISSOCIATION OF RYDBERG STATES OF CO INVESTIGATED BY THE DETECTION OF C(^{3}P_{J}) FRAGMENT

\underline{TAKUMICHI SUTANI}, TAKAYUKI EBATA, AND NAOHIKO MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai, 980-77, Japan.

Carbon atom generation following the predissociation of Rydberg states of CO has been investigated ^{a}. The Rydberg states of CO were prepared by two-color double resonant excitation through the 3sσB^{1}Σ^{+} state and the atomic carbon fragment was detected by resonance-enhanced multiphoton ionization. The 4s ∼ 6s (v = 0), 5s (v = 1), 4p ∼ 5p (v = 0), 3dπ (v = 0, 1), 4dσ (v = 0), 5dσ (v = 0), 4f ∼ 5f (v = 0) Rydberg states and the W^{1}Π(v = 2, 3) state were found to generate ^{3}P carbon atom. Mechanism of the predissociation of these high Rydberg states will be discussed based on the atomic carbon fragment yield measurement and on Doppler line shape measurement.

aT. Ebata, T. Sutani, and N. Mikami, Chem. Phys. Lett. 240, 357 (1995).